OpenStudy (anonymous):

I HAVE TWO STUPID QUESTIONS 1. 4 pipes can fill a tank in 70 minutes. How long will the tank fill when 7 pipes are used. Why do we multiply 70 by 4 and divide it by 7 and not the other way round? ALSO 2. When we do that map and actual thingy and theres the ratio of 1:5000 and the actual value it 1.6 cm and the map is unknown. What do I do?

OpenStudy (insa):

The pipes and time are inversely proportional.

OpenStudy (anonymous):

What about the other one?

OpenStudy (insa):

Im not that sure about it... Sorry

OpenStudy (welshfella):

1 : 5000 is the scale of the map it means the actual distances in the real world are 5000 times the distances on the map

OpenStudy (welshfella):

so if the distance is 1.6 on the map then real distance is 5000* 1.6

OpenStudy (anonymous):

THANK YOU GUYS

OpenStudy (perl):

another way to do this d= r*t 4 pipes = r * 70 min 4/70 = r d/r = t 7/ (4/70)= t t = 40